OpenStudy (anonymous):

point C is located at (1,2) and point D is located at (-4,-2). Find the x value for the point that is 1/4 the distance from point C to point D.

OpenStudy (anonymous):

let g(t)=(1 - 5 t, 2 - 4 t) t between 0 and 1. Notice that g(0)=(1,2)=C g(1)=(-4,-2)=D We need to compute g(1/5)=(0,6/5) This is your point. Why?

OpenStudy (anonymous):

Let us call A=(0,6/5). Compute the distance from A and from C and compare them \[ d_C=\frac{\sqrt{41}}{5}\\ d_D=\frac{4 \sqrt{41}}{5}\\ \frac {d_C}{d_A}=\frac 1 4 \]

OpenStudy (anonymous):

Another way is to compute \[ \frac 1 5 D + \frac 45 C=\frac 1 5(-4,-2) + \frac 45 (1, 2)=\\ \left(-\frac 4 5 + \frac 4 5, -\frac 2 5+ \frac 85 \right)=\left(0, \frac 6 5 \right) \]